OpenStudy (anonymous):

Find the derivative of y= (1+sinx) / (1-sin2x)

OpenStudy (anonymous):

dont know what i 've did wrong... i'm close to the answer..

OpenStudy (shayaan_mustafa):

HI cmkc109 :) first we have to use quotient rule. did you use that?

OpenStudy (anonymous):

yes. I used that

OpenStudy (anonymous):

answer given is cosx- cosxsin2x + 2cos2x + 2sinx cos / (1-sin2x)^2 but i only got cosx- cosxsin2x + 2cos2x / (1-sin2x)^2

OpenStudy (anonymous):

different ans than my ans key has.. i guess the one u have simplifies more..dont really get it lol.. never know u can find derivative in wolframalpha

OpenStudy (anonymous):

\[\frac{\cos (x)}{1-\sin (2 x)}+\frac{2 (\sin (x)+1) \cos (2 x)}{(1-\sin (2 x))^2} \]simplified is:\[\frac{-2 \sin ^3(x)+\cos (x)+2 \cos (2 x)}{(\sin (2 x)-1)^2} \]